OpenStudy (anonymous):

Write the equation of the parabola in vertex form. vertex (–4, 3), point (4, 131)

OpenStudy (anonymous):

there are two possibilities here... 4a(y-k) = (x-h)^2 or 4a(x-h) = (y-k)^2 where (h,k) is the vertex and a is the distance from the vertex to focus. to find a, stick in what you know from the given: h=-4 k=3 x=4 y=131 and solve for a in BOTH cases since it was not specified which type of parabola you have.